OpenStudy (anonymous):

A standard showerhead in Andrew's house dispenses 8 gallons of water per minute. Andrew changed this showerhead to an energy-saving one. The equation shows the amount of water dispensed, y, as a function of the number of minutes, x, for the new showerhead. y = 3x How much water does Andrew save each day with the change in showerheads if he uses the shower for 8 minutes each day? (1 point) 5 gallons 16 gallons 24 gallons 40 gallons

OpenStudy (anonymous):

@Directrix here's one

Directrix (directrix):

Standard Shower Head: 8 gallons of water per minute. y = 8 x New Shower Head: 3 gallons of water per minute. y = 3x Andrew showers 8 minutes per day. y = 8* 8 = 64 on the old shower head Y = 3*8 = 24 on the new shower head @Samirah16 How much water is Andrew saving per day based on this information?

Directrix (directrix):

Post what you think, okay?

OpenStudy (anonymous):

Do i subtract because i did and got 40

Directrix (directrix):

Correct. 40 gallons of water saved

OpenStudy (anonymous):

Thank you sooo much

OpenStudy (anonymous):

Would you like another metal?

Directrix (directrix):

You are welcome.